# Do Whitepaper Claims Predict Market Behavior? Evidence from Cryptocurrency Factor Analysis

### v1

@@ Abstract
Cryptocurrency projects articulate their value propositions through whitepapers, making foundational claims about functionality, use cases, and technical capabilities. This study investigates whether these narrative claims align with empirically observed market behavior. We construct a novel pipeline combining natural language processing (NLP) with tensor decomposition to compare three representational spaces: (1) a claims matrix derived from zero-shot classification of 24 whitepapers across 10 semantic categories using BART-MNLI, (2) a market statistics matrix capturing 7 financial metrics for 49 cryptocurrency assets over two years of hourly data (17,543 timestamps), and (3) latent factors extracted via CP tensor decomposition (rank 2, explaining 92.45% of variance). Using Procrustes rotation and Tucker’s congruence coefficient ($\phi$), we test alignment between narrative and market spaces across 23 common entities.
Results indicate weak alignment: claims–statistics ($\phi=0.341$, $p=0.332$), claims–factors ($\phi=0.077$, $p=0.747$; a less conservative non-padded estimate yields $\phi\approx 0.31$ in matched dimensions, see Appendix E), and statistics–factors ($\phi=0.197$, $p<0.001$). Critically, the statistics–factors significance ($p<0.001$) validates our methodology: since both matrices derive from the same market data, significant alignment confirms the Procrustes pipeline detects relationships when present, establishing that the narrative null result represents observed weak alignment rather than measurement failure. Inter-model validation using DeBERTa-v3 yields 32% exact agreement (Cohen’s $\kappa=0.14$) but 67% relaxed (top-3) agreement, indicating models capture similar semantic neighborhoods with different decision boundaries. Cross-sectional analysis reveals NEAR, MKR, and ATOM exhibit modest positive alignment contributions, while ENS, UNI, and Bitcoin show the largest divergence. Excluding Bitcoin ($n=22$) yields $\phi=0.081$ for claims–factors, confirming the null result is not driven by Bitcoin’s dominant market position.
We interpret these findings as consistent with weak alignment between whitepaper narratives and realized market factor structure. Limited statistical power ($n=23$) precludes distinguishing weak alignment from no alignment, but we can confidently reject strong alignment ($\phi\geq 0.70$). Implications for narrative economics, market efficiency, and investment analysis are discussed.
@@ 1 Introduction
Our findings reveal weak alignment across comparisons ($\phi<0.35$), with one notable exception: the statistics–factors comparison achieves statistical significance ($p<0.001$) despite weak magnitude, indicating systematic correspondence between market metrics and tensor-derived factors that narrative claims fail to capture. Infrastructure tokens (NEAR, MKR, ATOM) show modest positive alignment, while ENS, UNI, and Bitcoin exhibit the largest narrative-market divergence. This result is robust to temporal variation, subsample perturbation, decomposition method (CP vs Tucker), and rank selection.
@@ 2.1 Cryptocurrency Narratives and Sentiment
Research on cryptocurrency narratives spans social media analysis, whitepaper studies, and sentiment measurement. Chen et al. (2019) show that Twitter sentiment predicts Bitcoin returns at short horizons, while Ante (2023) find that Elon Musk’s tweets generate significant abnormal returns for mentioned cryptocurrencies. Haykir and Yağlı (2022) document speculative bubbles driven by herding behavior across crypto assets. Liu and Tsyvinski (2021) establish that cryptocurrency returns exhibit momentum, size effects, and exposure to market-wide factors distinct from traditional assets.
@@
Whitepaper analysis has received comparatively less attention. Howell et al. (2020) examine ICO whitepaper quality as a signal of project legitimacy, finding that technical depth correlates with fundraising success. Fisch (2019) show that whitepaper informativeness predicts ICO outcomes, and Florysiak and Schandlbauer (2022) find that analyst assessments of whitepaper quality correlate with funding success. However, these studies focus on cross-sectional prediction at issuance rather than ongoing alignment between narrative and market behavior. Indeed, Suriano et al. (2025) find that clustering cryptocurrencies by whitepaper content does not yield significant differences in time series dynamics, suggesting narratives may matter for initial fundraising but not long-term price behavior. Recent high-profile failures illustrate the gap between whitepaper claims and realized outcomes: Briola et al. (2022) analyze the Terra-Luna collapse, demonstrating how algorithmic stablecoin mechanisms failed despite elaborate whitepaper specifications, while Vidal-Tomás et al. (2023) document how centralized exchange fragility propagates through cryptocurrency markets.
Narrative Economics Framework. Shiller (2017) provides the theoretical foundation for studying how narratives shape economic outcomes, emphasizing how stories spread virally and influence collective behavior. Applied to cryptocurrency, this framework suggests that project narratives—embodied in whitepapers—should influence investor beliefs and thus market prices. Our study tests a related hypothesis: whether project narratives exhibit structural alignment with market factor structure. This differs from Shiller’s focus on narrative contagion dynamics; we examine whether the content of narratives corresponds to realized market behavior, finding limited support for such correspondence. This complements information-theoretic approaches (Keskin and Aste, 2020) that find nonlinear causality between social sentiment and cryptocurrency returns.
@@
Natural Language Processing in Finance. The application of NLP to financial text has grown substantially, with transformer-based models enabling sophisticated document analysis. Zero-shot classification, as employed here, allows categorization without domain-specific training data (Lewis et al., 2020). Bartolucci et al. (2020) demonstrate that sentiment extracted from developer communications (GitHub) predicts cryptocurrency price movements, validating NLP approaches in this domain. However, the trade-off between scalability and domain expertise remains an open question. Our use of BART-MNLI represents a middle ground—leveraging powerful pre-trained representations while acknowledging potential limitations in cryptocurrency-specific semantics.
@@ 2.2 Factor Models in Cryptocurrency
Traditional asset pricing employs factor models to explain cross-sectional return variation. Fama and French (1993) established the theoretical foundation for factor-based explanations of cross-sectional returns. Livan et al. (2011) demonstrate how random matrix theory can distinguish signal from noise in financial correlation matrices—a perspective we extend to narrative-factor comparisons. Caccioli et al. (2018) provide a comprehensive review of network-based approaches to financial systemic risk, complementing factor-based perspectives with topological analysis. In cryptocurrency, factor model development remains nascent but growing.
Liu and Tsyvinski (2021) propose a three-factor model for cryptocurrency returns: market, size, and momentum. They find that these factors explain substantial cross-sectional variation, analogous to the Fama-French factors for equities. Our tensor decomposition implicitly captures similar factors—Factor 1 (dominated by Bitcoin) resembles their market factor, while Factor 2 may capture size or sector effects.
Recent work has explored additional factors specific to cryptocurrency: network effects (measured by active addresses), tokenomics (supply schedule, staking yields), and on-chain activity (transaction volume, developer commits). These factors may mediate narrative-market relationships in ways our current analysis does not capture.
@@ 2.3 Tensor Methods in Finance
Tensor decomposition provides a natural framework for analyzing multi-way financial data. Kolda and Bader (2009) provide a comprehensive review of tensor decomposition methods, establishing the theoretical foundations for CP and Tucker decomposition. In finance, Chen et al. (2022) develop factor models for high-dimensional tensor time series, establishing the statistical foundations for tensor factor analysis. Fan et al. (2013) introduce the POET estimator for high-dimensional covariance estimation with factor structure.
@@ 2.4 Factor Comparison Methods
Comparing factor structures across studies or datasets requires methods that account for rotational indeterminacy. Procrustes rotation (Schönemann, 1966) finds the optimal orthogonal transformation aligning one factor matrix to another. Tucker’s congruence coefficient ($\phi$) then measures similarity between aligned factors (Tucker, 1951; Lorenzo-Seva and ten Berge, 2006).
Lorenzo-Seva and ten Berge (2006) establish interpretation thresholds: $|\phi|\geq 0.95$ indicates factor equivalence, $|\phi|\geq 0.85$ indicates fair similarity, $|\phi|\geq 0.65$ indicates moderate similarity, and $|\phi|<0.65$ indicates weak or no similarity. These thresholds guide our interpretation of narrative-market alignment.
@@ 3.1 Market Data
Assets (whitepapers) 24
Assets (common intersection) 23
@@ 3.2 Whitepaper Corpus
We collect whitepapers for 24 assets where official foundational documents are publicly available: AAVE, ADA, ALGO, AR, ARB, ATOM, AVAX, BTC, COMP, DOT, ENS, ETH, FIL, GRT, ICP, LINK, MKR, NEAR, SC, SOL, STORJ, UNI, XMR, and ZEC. Documents include original whitepapers, consensus papers, protocol specifications, and technical documentation. PDF text is extracted using PyPDF2 with sentence-level tokenization; for assets without extractable PDFs, we use official documentation in markdown format. Table 2 summarizes corpus statistics for selected assets.
@@
The intersection of whitepaper and market data yields 23 common assets for alignment analysis (AR lacks sufficient market data coverage in our sample).
@@ 4.1 Tensor Construction
With a single venue, the effective structure is 3-way: $\mathcal{X}\in\mathbb{R}^{T\times A\times F}$. Each entry $x_{taf}$ represents the value of feature $f$ for asset $a$ at time $t$.
@@ 4.2.3 Rank Selection
With target $\text{EV}\geq 0.90$, we obtain $R=2$ (EV = 92.45%). [^1] Standard PCA on mode-2 (asset) matricization yields nearly identical alignment ($\phi=0.060$, vs CP $\phi=0.060$), confirming that preserving tensor structure does not alter our findings. We retain the tensor framework for interpretability of multi-way temporal dynamics and consistency with recent financial tensor methods (Chen et al., 2022).
@@ 4.4 Market Statistics
We compute seven z-normalized summary statistics for each asset:
@@ 4.7.1 Power Considerations
With $n=23$ common entities, statistical power to detect alignment is limited. Monte Carlo simulation (500 iterations per effect size, 200 permutations each) reveals approximate power at $\alpha=0.05$:
@@ 4.7.3 Bootstrap Confidence Intervals
We construct 95% CIs via percentile bootstrap ($B=1000$ resamples). However, bootstrap resampling with replacement on small samples ($n=23$) exhibits known pathologies when combined with Procrustes-based alignment: duplicate entities in resampled data artificially inflate $\phi$ by increasing effective weights on well-aligned pairs. Our bootstrap distributions show substantial upward bias (bootstrap mean exceeds point estimate by 29% for claims–statistics, 37% for claims–factors), with moderate right-skewness (skewness $\approx 0.45$). Consequently, percentile CIs may be conservative for upper bounds but unreliable for lower bounds. Given this bias, reported confidence intervals should be interpreted as indicative rather than precise. We report these intervals for completeness while acknowledging their limitations.

### v2

@@ Abstract
Cryptocurrency projects articulate their value propositions through whitepapers, making foundational claims about functionality, use cases, and technical capabilities. This study investigates whether these narrative claims align with empirically observed market behavior. We construct a novel pipeline combining natural language processing (NLP) with tensor decomposition to compare three representational spaces: (1) a claims matrix derived from zero-shot classification of 38 whitepapers across 10 semantic categories using BART-MNLI, (2) a market statistics matrix capturing 7 financial metrics for 49 cryptocurrency assets over two years of hourly data (17,543 timestamps), and (3) latent factors extracted via CP tensor decomposition (rank 2, explaining 92.45% of variance). Using Procrustes rotation and Tucker’s congruence coefficient ($\phi$), we test alignment between narrative and market spaces across 37 common entities.
Results indicate weak alignment: claims–statistics ($\phi=0.246$, $p=0.339$), claims–factors ($\phi=0.058$, $p=0.751$), and statistics–factors ($\phi=0.174$, $p<0.001$). Critically, the statistics–factors significance ($p<0.001$) validates our methodology: since both matrices derive from the same market data, significant alignment confirms the Procrustes pipeline detects relationships when present, establishing that the narrative null result represents observed weak alignment rather than measurement failure. Inter-model validation using DeBERTa-v3 yields 32% exact agreement (Cohen’s $\kappa=0.14$) but 67% relaxed (top-3) agreement, indicating models capture similar semantic neighborhoods with different decision boundaries. Cross-sectional analysis reveals XMR, CRV, YFI, and SOL exhibit positive alignment contributions, while RPL, HBAR, AAVE, and SUSHI show the largest divergence.
We interpret these findings as consistent with weak alignment between whitepaper narratives and realized market factor structure. Limited statistical power ($n=37$) precludes distinguishing weak alignment from no alignment, but we can confidently reject strong alignment ($\phi\geq 0.70$). Implications for narrative economics, market efficiency, and investment analysis are discussed.
@@ 1 Introduction
Our findings reveal weak alignment across comparisons ($\phi<0.25$), with one notable exception: the statistics–factors comparison achieves statistical significance ($p<0.001$) despite weak magnitude, indicating systematic correspondence between market metrics and tensor-derived factors that narrative claims fail to capture. Specialized tokens (XMR, CRV, YFI, SOL) show positive alignment contributions, while DeFi infrastructure tokens (RPL, HBAR, AAVE, SUSHI) exhibit the largest narrative-market divergence. This result is robust to temporal variation, subsample perturbation, decomposition method (CP vs Tucker), and rank selection.
@@ 2.1 Cryptocurrency Narratives and Sentiment
Research on cryptocurrency narratives spans social media analysis, whitepaper studies, and sentiment measurement. Chen et al. (2019) show that Twitter sentiment predicts Bitcoin returns at short horizons, while Ante (2023) find that Elon Musk’s tweets generate significant abnormal returns for mentioned cryptocurrencies. Haykir and Yağlı (2022) document speculative bubbles driven by narrative contagion across crypto assets. Liu and Tsyvinski (2021) establish that cryptocurrency returns exhibit momentum, size effects, and exposure to market-wide factors distinct from traditional assets.
Whitepaper analysis has received growing attention as a signal of project quality. Howell et al. (2020) examine ICO whitepaper quality as a signal of project legitimacy, finding that technical depth correlates with fundraising success. Fisch (2019) show that whitepaper informativeness predicts ICO outcomes. Adhami et al. (2018) provide the first comprehensive empirical description of ICO determinants, finding that code availability, token presale, and utility design predict success. More recent work has applied sophisticated NLP to whitepaper content: Thewissen et al. (2022) use topic modeling on 5,210 whitepapers to show that technical feature topics predict success while whitepaper informativeness diminishes post-listing. Critically, Momtaz (2021) demonstrates that issuers systematically exaggerate whitepaper claims—exaggeration raises funds short-term but causes token depreciation and platform failure long-term. Samieifar and Baur (2021) find that whitepaper length and complexity correlate positively with funds raised, supporting signaling interpretations. Florysiak and Schandlbauer (2022) show that while expert ratings initially “jam” whitepaper signals, post-listing returns are better predicted by whitepaper content than analyst assessments.
@@
However, these studies focus on cross-sectional prediction at issuance rather than ongoing alignment between narrative and market behavior. Indeed, Suriano et al. (2025) find that clustering cryptocurrencies by whitepaper content does not yield significant differences in time series dynamics, suggesting narratives may matter for initial fundraising but not long-term price behavior. Recent high-profile failures illustrate the gap between whitepaper claims and realized outcomes: Briola et al. (2022) analyze the Terra-Luna collapse, demonstrating how algorithmic stablecoin mechanisms failed despite elaborate whitepaper specifications, while Vidal-Tomás et al. (2023) document how centralized exchange fragility propagates through cryptocurrency markets. Our work extends this by testing whether claims align with ongoing market factor structure.
Narrative Economics Framework. Shiller (2017) provides the theoretical foundation for studying how narratives shape economic outcomes. This builds on foundational work in behavioral finance: Baker and Wurgler (2006) construct the canonical investor sentiment index, demonstrating that sentiment-driven mispricing is strongest for speculative, hard-to-value stocks—a category that includes most cryptocurrencies. Baker and Wurgler (2007) survey evidence on how sentiment affects both aggregate returns and cross-sectional pricing. Tetlock (2007) shows that media pessimism predicts market downturns and high trading volume, establishing the textual analysis framework for investor sentiment. Applied to cryptocurrency, this framework suggests that project narratives—embodied in whitepapers—should influence investor beliefs and thus market prices. Our study tests this prediction empirically, finding weak support for narrative-market coupling. This complements information-theoretic approaches (Keskin and Aste, 2020) that find nonlinear causality between social sentiment and cryptocurrency returns.
@@
Natural Language Processing in Finance. The application of NLP to financial text has grown substantially, with transformer-based models enabling sophisticated document analysis (Loughran and McDonald, 2020). Loughran and McDonald (2011) demonstrate that general sentiment dictionaries misclassify financial text, motivating domain-specific approaches. Kearney and Liu (2014) provide a comprehensive methodological survey of textual sentiment methods and models. Domain-specific pre-training has proven effective: Araci (2019) introduce FinBERT, outperforming Loughran-McDonald dictionaries on financial sentiment, while Huang et al. (2020) show similar gains on analyst reports. Mishev et al. (2020) benchmark lexicons against transformers (BERT, FinBERT) across financial datasets, finding transformer superiority for nuanced sentiment. Zero-shot classification, as employed here, allows categorization without domain-specific training data (Lewis et al., 2020). Bartolucci et al. (2020) demonstrate that sentiment extracted from developer communications (GitHub) predicts cryptocurrency price movements, validating NLP approaches in this domain. Our use of BART-MNLI represents a middle ground—leveraging powerful pre-trained representations while acknowledging potential limitations in cryptocurrency-specific semantics.
@@ 2.2 Factor Models in Cryptocurrency
Traditional asset pricing employs factor models to explain cross-sectional return variation. Fama (1970) established the theoretical foundation for efficient markets and factor-based returns. Fama and French (1993) introduced the three-factor model for equities; analogous developments in cryptocurrency have emerged more recently. Livan et al. (2011) demonstrate how random matrix theory can distinguish signal from noise in financial correlation matrices—a perspective we extend to narrative-factor comparisons. Caccioli et al. (2018) provide a comprehensive review of network-based approaches to financial systemic risk, complementing factor-based perspectives with topological analysis.
Liu et al. (2019) establish the foundational three-factor model for cryptocurrency returns: market, size, and momentum. Liu and Tsyvinski (2021) extend this work, finding that these factors explain substantial cross-sectional variation analogous to Fama-French factors for equities. Bianchi and Babiak (2021) apply Instrumented PCA (IPCA) to show that time-varying factor loadings outperform observable risk factors. Dobrynskaya (2020) extends crypto CAPM to include downside beta as a fourth factor, finding significant cross-sectional premia across 1,700 coins. Bhambhwani et al. (2019) introduce blockchain-native factors—computing power, network size—as procyclical pricing factors with positive risk premia.
Our tensor decomposition implicitly captures similar factors—Factor 1 (dominated by Bitcoin) resembles the market factor, while Factor 2 may capture size or sector effects. These systematic factors may dominate any narrative-based signal, explaining why whitepaper claims fail to predict factor structure.
@@ 2.3 Tensor Methods in Finance
Tensor decomposition provides a natural framework for analyzing multi-way financial data. Kolda and Bader (2009) provide a comprehensive review of tensor decomposition methods, establishing the theoretical foundations for CP and Tucker decomposition. Recent work has applied these methods to financial time series: Chen et al. (2022) develop tensor factor models for high-dimensional time series, introducing TIPUP/TOPUP estimators with explicit applications to economics and finance. Wang et al. (2022) apply Tucker decomposition to high-dimensional vector autoregression, demonstrating improved performance over matrix-based approaches for multivariate time series. Han et al. (2024) develop CP factor models for dynamic tensors, providing uncorrelated latent factors directly applicable to asset pricing. Fan et al. (2013) introduce the POET estimator for high-dimensional covariance estimation with factor structure.
@@ 2.4 Factor Comparison Methods
Comparing factor structures across studies or datasets requires methods that account for rotational indeterminacy. Procrustes rotation (Schönemann, 1966) finds the optimal orthogonal transformation aligning one factor matrix to another. Brokken (1983) develops orthogonal Procrustes rotation that directly maximizes congruence, the approach we employ. Tucker’s congruence coefficient ($\phi$) then measures similarity between aligned factors (Tucker, 1951; Lorenzo-Seva and ten Berge, 2006).
Korth and Tucker (1975) establish null distributions for congruence coefficients from simulated data, essential for understanding what constitutes statistically meaningful alignment. Paunonen (1997) examines the distribution of factor congruence under chance conditions after Procrustes rotation, informing our significance testing. Lorenzo-Seva and ten Berge (2006) establish interpretation thresholds: $|\phi|\geq 0.95$ indicates factor equivalence, $|\phi|\geq 0.85$ indicates fair similarity, $|\phi|\geq 0.65$ indicates moderate similarity, and $|\phi|<0.65$ indicates weak or no similarity. These thresholds guide our interpretation of narrative-market alignment.
@@ 3.1 Market Data
Assets (whitepapers) 38
Assets (common intersection) 37
@@ 3.2 Whitepaper Corpus
We collect whitepapers for 38 assets where official foundational documents are publicly available: AAVE, ADA, ALGO, AR, ARB, ATOM, AVAX, BTC, COMP, DOT, ENS, ETH, FIL, GRT, ICP, LINK, MKR, NEAR, SC, SOL, STORJ, UNI, XMR, and ZEC. Documents include original whitepapers, consensus papers, protocol specifications, and technical documentation. PDF text is extracted using PyPDF2 with sentence-level tokenization; for assets without extractable PDFs, we use official documentation in markdown format. Table 2 summarizes corpus statistics for selected assets.
@@
Table 3 summarizes data availability across the pipeline. The intersection of whitepaper ($n=38$) and market data ($n=49$) yields 37 common assets for alignment analysis. AR is excluded due to insufficient market data coverage.
Table 3: Data Flow and Asset Coverage
Data Source Assets Notes
Whitepaper corpus 38 Technical documents
Market data (Binance) 49 2-year hourly OHLCV
Tensor factors (CP) 49 Rank-2 decomposition
NLP $\cap$ Market 37 Primary analysis sample
Excluded 1 AR (market data gap)
Market-only 12 No whitepaper available
@@ 4.1 Tensor Construction
With a single venue, the effective structure is 3-way: $\mathcal{X}\in\mathbb{R}^{T\times A\times F}$. Each entry $x_{taf}$ represents the value of feature $f$ for asset $a$ at time $t$. Prior to decomposition, we z-normalize each feature slice across both assets and time (i.e., each $\mathcal{X}_{::f}$ has zero mean and unit variance), ensuring that scale differences across OHLCV features do not dominate the factor structure.
@@ 4.2.3 Rank Selection
With target $\text{EV}\geq 0.90$, we obtain $R=2$ (EV = 92.45%). [^1] Standard PCA on mode-2 (asset) matricization yields nearly identical alignment ($\phi=0.058$, vs CP $\phi=0.058$), confirming that preserving tensor structure does not alter our findings. We retain the tensor framework for interpretability of multi-way temporal dynamics and consistency with recent financial tensor methods (Chen et al., 2022).
@@ 4.3.3 Model Validation
Multi-Method Validation.
To further assess classification robustness, we implement three independent methods with distinct inductive biases: (1) BART-MNLI (NLI-based entailment), (2) sentence embeddings using all-mpnet-base-v2 (Reimers and Gurevych, 2019) with cosine similarity to category descriptions, and (3) Ministral-3 3B, a local language model via structured JSON prompting. Table 5 reports pairwise correlations across the 38-asset expanded corpus.
Table 5: Multi-Method Classification Agreement
Method Pair Pearson $r$ Spearman $\rho$
BART-NLI vs Embedding 0.103 0.111
BART-NLI vs LLM 0.413 0.457
Embedding vs LLM 0.392 0.384
Mean pairwise 0.302 0.317
The LLM-based classifier exhibits moderate correlation with both other methods ($r\approx 0.4$), while BART and embedding methods show weaker agreement ($r=0.10$), suggesting distinct inductive biases. Discretized Fleiss’ Kappa ($\kappa=0.045$) indicates slight but positive inter-rater agreement above chance.
Notably, per-category agreement varies substantially (see Appendix G for full heatmap): categories with clear linguistic markers show strong convergence (medium_of_exchange $\bar{r}=0.65$, DeFi $\bar{r}=0.63$, privacy $\bar{r}=0.54$), while abstract concepts show weaker agreement (smart_contracts $\bar{r}=0.24$, store_of_value $\bar{r}=0.38$). This pattern suggests that some functional categories are more robustly identifiable across methods than others—a finding with implications for taxonomy design in cryptocurrency NLP research. Figure 1 visualizes these cross-method patterns.
Figure 1: Multi-method classification comparison across 10 semantic categories. Bar heights represent mean classification scores for BART-NLI (primary), sentence embeddings, and Ministral-3 LLM. Categories with high cross-method agreement (medium_of_exchange, smart_contracts, DeFi) show consistent relative rankings; categories with weak agreement (store_of_value, governance) exhibit larger inter-method variance.
Data Quality Filtering.
Seven assets (ALGO, AXS, BAND, DOT, RPL, SUSHI, YFI) contain fewer than 10 text chunks, yielding unreliable classification estimates. Excluding these low-data assets ($n=31$ reliable), the embedding–LLM correlation improves substantially ($r=0.52$), suggesting that classification disagreement partially reflects data sparsity rather than fundamental method divergence.
@@ 4.4 Market Statistics
We compute seven summary statistics for each asset, then z-normalize cross-sectionally (across assets) to ensure comparability:
@@ 4.7.1 Power Considerations
With $n=37$ common entities, statistical power to detect alignment is limited. Monte Carlo simulation (500 iterations per effect size, 200 permutations each) reveals approximate power at $\alpha=0.05$:
@@ 4.7.3 Bootstrap Confidence Intervals
We construct 95% CIs via percentile bootstrap ($B=1000$ resamples). However, bootstrap resampling with replacement on small samples ($n=37$) exhibits known pathologies when combined with Procrustes-based alignment: duplicate entities in resampled data artificially inflate $\phi$ by increasing effective weights on well-aligned pairs. Our bootstrap distributions show substantial upward bias (bootstrap mean exceeds point estimate by 29% for claims–statistics, 37% for claims–factors), with moderate right-skewness (skewness $\approx 0.45$). Consequently, percentile CIs may be conservative for upper bounds but unreliable for lower bounds. Given this bias, reported confidence intervals should be interpreted as indicative rather than precise. We report these intervals for completeness while acknowledging their limitations.
4.7.4 Reproducibility
All stochastic procedures use fixed random seeds (seed = 42 for CP-ALS initialization, tensor operations, and permutation tests) to ensure reproducibility. Permutation tests use $B=1000$ iterations; bootstrap procedures use $B=1000$ resamples. Implementation uses Python 3.11+ with NumPy, SciPy, TensorLy (CP-ALS), scikit-learn (CCA, PLS), and Hugging Face Transformers (BART-MNLI). Code and data are available at [repository URL].